Felix Enterprise is a small retail business which specialises in children’s clothing. The company is considering buying a cash register software so that it can effectively deal with its retail sales. The software package costs GHS 750,000 and will be depreciated down to zero using straight line method over its five-year economic life. The marketing department of the company predicts sales will be GHS 600,000 per year for the next three years, after which the market will cease to exist. Cost of goods sold and operating expenses are predicted to be 25 percent of sales. After three years, the software can be sold for GHS 40,000. Felix Enterprise also needs to add net working capital of GHS 25,000 immediately. The additional net working capital will be recovered in full at the end of the project’s life. The Finance department of the company has indicated that the company’s marginal tax rate is 35% and the project has a beta of 1.25. The current risk free rate is 14.75% and the market return for such project is 21%. Using the NPV method, determine whether Felix Enterprise should purchase this software.
Answer #1

Solution:

From the above question, it is evident that the project's life is 3 years, as the sales are projected for 3 years from the asset which would be sold in the third year.

Step 1: Calculation of Tax saving on Net loss on assets

Calculation for Depreciation ( Straight line ) for 3 years

(Assets Cost/ Economic life )* No of years = (GHS 750,000/5)*3 = GHS 450,000

Written Down Value at the end of 3rd year = GHS 750,000- GHS 450,000= GHS 300,000

This software was sold for GHS 40,000

Net loss on software =GHS (300,000-40,000) = GHS 260,000

Tax shield on Net loss on software (inflow) = 260,000*35%= GHS 91,000

Step 2: Calculation of Discounting Rate

Using Capital Asset Pricing Model ( CAPM), Expected Return of Project = Interest Free + Beta ( Market Return- Interest Free)

Interest Free = 14.75%

Beta = 1.25

Market Return= 21%

Expected Return of the Project = 14.75% + 1.25 ( 21%-14.75%) = 14.75+7.81= 22.56%

Step 3 : Terminal Value Cash Flows ( at the end of Year 3)

Present Value factor for the rate of 22.56% at the end of year 3 is 0.5431

Particulars Amount ( In GHS)

a)Tax shield on Loss of Assets ( as computed in Step 1)

91,000
b) Cash Inflow from Sale of Assets 40,000
c) Net Cash Inflow (a+b) 131,000
d) Discounting Factor 0.5431
e) Discounted Value (c*d) 71,158.21
f) Additional Recovery of Working Capital 25,000
g) Discounted Value of Working Capital (f*d) 13,577.56
h) Total Cash Inflow ( e + g) 84,735.71

Step 4 : Calculation of Free Cash Flows (FCF)

Years Sales ( In GHS) COGS & Operating Expenses ( In GHS) @ 25% of Sales Depreciation ( In GHS) Earning Before Taxes (In GHS) Tax Rate @ 35% Earning After Taxes (EAT in GHS)

FCF (EAT+Depreciation)

Discounting factor @22.56% Discounted Cash Flows ( GHS)
1 600,000 150,000 150,000 300,000 105,000 195,000 345,000 0.8159 281485.50
2 600,000 150,000 150,000 300,000 105,000 195,000 345,000 0.6657 229,666.50
3 600,000 150,000 150,000 300,000 105,000 195,000 345,000 0.5431 187,369.50
Discounted Cash Inflow 698,521.50

Total Cash Inflow = Discounted Cash Inflow + Terminal Value Cash Flows = GHS 698,522 +84,736= GHS 783,258

Step 5 : Calculation of Cash Outflow

Initial Cash outlay at Year 0 = GHS 750,000

Add: Additional Working Capital immediately ( means at Year 0) = GHS 25,000

Net Cash outflow ( In GHS) =GHS 775,000

Step 6 : Net Present Value

NPV = Cash Inflow - Cash Outflow = ( GHS 783,258-GHS 775,000) = GHS 8,258

Recommendation:

  • The Felix Enterprise should purchase this software as the net Present Value is GHS 8,258